HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

UV SPECTRUM AND PHOTODECOMPOSITION OF PEROXYNITROUS ACID (HOONO)

\underline{W. CHEBBI} , GSMA, Groupe de Spectrométrie Moléculaire et Atmosphérique, UMR CNRS 7331, University of Reims Champagne-Ardenne, 51100 Reims, France;

LSAMA, Laboratoire de Spectroscopie Atomique, Moléculaire et Applications,

Department of Physics, University Tunis - El Manar, 1060 Tunis, Tunisia; N. DERBEL, LSAMA, Laboratoire de Spectroscopie Atomique, Moléculaire et Applications, Department of Physics, University Tunis - El Manar, 1060 Tunis,

Tunisia; T. COURS, A. ALIJAH, GSMA, Groupe de Spectrométrie Moléculaire et Atmosphérique, UMR CNRS 7331, University of Reims Champagne-Ardenne, 51100

Reims, France

Peroxynitrous acid HOONO plays an important role in atmospheric chemistry. It is formed as a minor product by the chain-terminating reaction of NO_{2} and OH, the main product being nitric acid, HONO_{2}. Peroxynitrous acid was first observed in an Argon matrix^{1}. Its UV spectrum was recorded soon after^{2}. In contrast to nitric acid, peroxynitrous acid is an unstable species and has not been seen in the gas phase at ambient pressure and temperature because of its short lifetime of a few seconds^{3}. HOONO disintegrates readily through thermolysis and photodecomposition. These processes are not well understood and are investigated in the present work.

To this end, the UV spectrum was computed using both DFT and ab initio methods. Upon UV excitation, the molecule will break up to yield either HO_{2} + NO or NO_{2} + OH. The decay process was studied in detail by non-adiabatic dynamics on the lowest four electronic potential energy surfaces using the NEWTON-X package^{4}, interfaced with the MOLPRO and GAUSSIAN programs. From a statistical analysis of a large number of trajectories, we were able to obtain the fragment distributions for the two conformers cis-cis and trans-perp. Our analysis revealed that the cis-cis conformer produces 95% of HOO+NO and 2% of HO+ONO, whereas the trans-perp conformer yields HOO+NO with a 100% yield. The tropospheric photolysis rate constant was then obtained from the computed UV cross-section data and the actinic flux.

1Cheng B.-M., Lee J.-W. and Lee Y.-P., J. Phys. Chem, 95, 2814-2817, (1991).

2W.J Lo, Y.P Lee, Chem. Phys. Lett, 229, 357, (1994).

3Abida, O., Mielke, L.H., Osthoff, H.D., Chem. Phys. Lett, 511, 187-192, (2011).

4M. Barbatti, Wiley Interdiscip. Rev.: Comput. Mol. Sci, 1, 620–633, (2011).

p-number: p045  Submitted on Wed, 07 Jun 2023 13:24:30 +0200